# The Sample-Communication Complexity Trade-off in Federated Q-Learning

### v1

@@ A.5.1 Proof of Lemma 2
Note that a similar relationship is also derived in Li et al. [2024], but needing to take care of the averaging over multiple agents, we present the entire arguments for completeness. We prove the claim using an induction over $t$. It is straightforward to note that the claim is true for $t=0$ and all agents $m\in\{1,2,\dots,M\}$. For the inductive step, we assume that the claim holds for $t-1$ for all clients. Using the induction hypothesis, we have the following relation between $V_{t-1}^{m}(1)$ and $\widehat{V}_{t-1}^{m}$:
@@ A.5.3 Proof of Lemma 4
Let $\{\mathscr{F}_{l}\}_{l=1}^{L}$ be a filtration such that $\mathscr{F}_{l}$ is the $\sigma$-algebra corresponding to $\{P_{k(j),b(j)}^{m}(1|1,1),P_{k(j),b(j)}^{m}(1|1,2)\}_{j=1}^{l}$. It is straightforward to note that $\{z_{l}\}_{l=1}^{L}$ is a martingale sequence adapted to the filtration {}$\{\mathscr{F}\}_{l=1}^{L}$. We will use the Freedman’s inequality [Freedman, 1975, Li et al., 2024] to obtain a high probability bound on $|\zeta_{t^{\prime},t}^{m}|$.
@@
$\geq\sum_{l=1}^{L}\left(\widetilde{\eta}_{k(l)}^{(t)}\right)^{2}\frac{\gamma^{2}}{B^{2}}\cdot 2p(1-p)\cdot\mathbb{E}\left[\left(\widehat{V}_{k(l)-1}^{m}\right)^{2}\right]$
@@ A.5.4 Proof of Lemma 5
$\leq\left(\sum_{k=1}^{T}\mathbb{E}\left[\left(\sum_{m=1}^{M}\widetilde{\eta}_{k}^{(T)}\cdot\frac{\gamma}{M}\cdot(\widehat{P}_{k}^{m}(1|1,1)-\widehat{P}_{k}^{m}(1|1,2))\widehat{V}_{k-1}^{m}\right)^{2}\right]\right)^{1/2}$
$\leq\left(\sum_{k=1}^{T}\left(\widetilde{\eta}_{k}^{(T)}\right)^{2}\cdot\frac{2\gamma^{2}p(1-p)}{BM^{2}}\cdot\sum_{m=1}^{M}\mathbb{E}\left[\left(\widehat{V}_{k-1}^{m}\right)^{2}\right]\right)^{1/2}$

### v2

@@ A.5.1 Proof of Lemma 2
Note that a similar relationship is also derived in Li2023QLMinimax, but needing to take care of the averaging over multiple agents, we present the entire arguments for completeness. We prove the claim using an induction over $t$. It is straightforward to note that the claim is true for $t=0$ and all agents $m\in\{1,2,\dots,M\}$. For the inductive step, we assume that the claim holds for $t-1$ for all clients. Using the induction hypothesis, we have the following relation between $V_{t-1}^{m}(1)$ and $\widehat{V}_{t-1}^{m}$:
@@ A.5.3 Proof of Lemma 4
Let $\{\mathscr{F}_{l}\}_{l=1}^{L}$ be a filtration such that $\mathscr{F}_{l}$ is the $\sigma$-algebra corresponding to $\{P_{k(j),b(j)}^{m}(1|1,1),P_{k(j),b(j)}^{m}(1|1,2)\}_{j=1}^{l}$. It is straightforward to note that $\{z_{l}\}_{l=1}^{L}$ is a martingale sequence adapted to the filtration {}$\{\mathscr{F}\}_{l=1}^{L}$. We will use the Freedman’s inequality (Freedman1975Martingale; Li2023QLMinimax) to obtain a high probability bound on $|\zeta_{t^{\prime},t}^{m}|$.
@@
$\geq\sum_{l=1}^{L}\left(\widetilde{\eta}_{k(l)}^{(t)}\right)^{2}\frac{\gamma^{2}}{B^{2}}\cdot 2p(1-p)\cdot\mathbb{E}\left[\left(\hat{V}_{k(l)-1}^{m}\right)^{2}\right]$
@@ A.5.4 Proof of Lemma 5
$\leq\left(\sum_{k=1}^{T}\mathbb{E}\left[\left(\sum_{m=1}^{M}\widetilde{\eta}_{k}^{(T)}\cdot\frac{\gamma}{M}\cdot(\widehat{P}_{k}^{m}(1|1,1)-\widehat{P}_{k}^{m}(1|1,2))\hat{V}_{k-1}^{m}\right)^{2}\right]\right)^{1/2}$
$\leq\left(\sum_{k=1}^{T}\left(\widetilde{\eta}_{k}^{(T)}\right)^{2}\cdot\frac{2\gamma^{2}p(1-p)}{BM^{2}}\cdot\sum_{m=1}^{M}\mathbb{E}\left[\left(\hat{V}_{k-1}^{m}\right)^{2}\right]\right)^{1/2}$